one number is 12 more than another. The sum of the smaller number and twice the larger number is 39. Find the larger number.
Answer by richard1234(7193)

Let the smaller number be a, and the larger number be a+12. Then,

a + 2(a+12) = 39

3a + 24 = 39

a = 5, a+12 = 17.